OpenStudy (anonymous):

Write the slope-intercept equation for the line that passes through (-1, -4) and (4, 11). Please show all of your work.

OpenStudy (anonymous):

\[y_{1} - y _{2} = m(x _{1}-x _{2})\] so \[-15 = m(-5)\] \[m = -15/-5\] \[m = 3\] Then sub in m with unknowns: \[(-4) -y = 3((-1)-x)\] therefore: \[y=3x+1\] Q.E.D

OpenStudy (anonymous):

very sorry, i mean: \[y = 3x-1\]

OpenStudy (anonymous):

thanx

OpenStudy (anonymous):

ooops sorry m=3,,

OpenStudy (anonymous):

to find the slope, use y2-y1=m(x2-x1) -4-11=m(-1-4) -15=m(-5) m=(-15)/(-5)=3 always use the slope formula y=mx + b here it passes through point (-1,-4)=(x,y) therefore use the formula y-y1=m(x-x1) y-(-4)=(3)(x-(-1)) y+4 =(3)(x+1) y+4=(3)x +(1/3) y=(3x)+(3)-4 y=3x-1 ans